OpenStudy (anonymous):

In order to build a scale model of the trail, the drawing is enlarged on the coordinate plane. If two corners of the trail are at point A (-2, 7) and point D (-10, -1), what is another point that could represent a corner of the trail?

OpenStudy (anonymous):

@mathslover do you think you can help?

OpenStudy (anonymous):

assuming the trail model is contained within a rectangle?

OpenStudy (anonymous):

The model is actually a triangle lol. Ill post a pic!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

This is it

OpenStudy (anonymous):

ok can you figure out the scale factor?

OpenStudy (anonymous):

@breariley1

OpenStudy (anonymous):

No.. Can you tell me how to do that? Im sorry i am terrible at geometry!

OpenStudy (anonymous):

start at point D and move up unit by unit, and count the number of units it takes to get on the same y coordinate (which is the same horizontal line) with point A

OpenStudy (anonymous):

let me know what number you get

OpenStudy (anonymous):

I got 10....

OpenStudy (anonymous):

close but its 8, you can also take the absolute value difference of the y coordinates, -1 -7 = -8, absolute value of -8 is 8, OR going the other way, 7 - (-1) = 7+1 = 8

OpenStudy (anonymous):

ok now do the same thing with the smaller paralellogram, start at point i and go up until same y coordinate as point F

OpenStudy (anonymous):

Where did the numbers 1&7 come from though ?

OpenStudy (anonymous):

the y coordinates of the 2 points

OpenStudy (anonymous):

points A and D

OpenStudy (anonymous):

The coordinates are -2 and 4... So

OpenStudy (anonymous):

the y coordinate of point A is 7 and the y coordinate of point D is -1

OpenStudy (anonymous):

2 and 2?

OpenStudy (anonymous):

what does 2 and 2 refer to?

OpenStudy (anonymous):

Points i and f

OpenStudy (anonymous):

ok lets start over, ignore what i said up top because that was if it was a rectangle, but its not

OpenStudy (anonymous):

Okay haha!

OpenStudy (anonymous):